OpenStudy (anonymous):

if you hit a home run 53% of the time. if you hit 10 balls, what is the probability that you hit 4 homeruns. A) 0.0789 B)0.179 C)0.0905 D)0.000851

OpenStudy (kropot72):

\[P(4\ home\ runs)=\left(\begin{matrix}10 \\ 4\end{matrix}\right)0.53^{4}0.47^{6}=210\times 0.53^{4}\times 0.47^{6}\]

OpenStudy (anonymous):

thanks so B

OpenStudy (kropot72):

Yes, B is correct.